Determine the amount of money, to the nearest dollar, you must invest at 5% per year, compounded annually, so that you will be a millionaire in 33 years.
Answer by Alan3354(69443)

Determine the amount of money, to the nearest dollar, you must invest at 5% per year, compounded annually, so that you will be a millionaire in 33 years.
---------
FV = PV*(1 + r)^t
1000000 = PV*(1 + 0.05)^33
PV = 1000000/(1.05)^33
Just calculator work.